# Linear Bandits with Partially Observable Features

### v1

@@ C.5 Proof of Theorem 5
with probability at least $1-3\delta/t$, by Theorem 4. Summing over $t$ gives,
@@
By resampling at most $\rho_{t}$ times, the probability of the event $\{a_{t}\neq\widehat{a}_{t}\}$ is
$\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)=$ $\sum_{m=1}^{\rho_{t}}\frac{p}{(K-1)\sqrt{t}}\left(1-\frac{p}{(K-1)\sqrt{t}}\right)^{m-1}$
$=$ $\frac{p}{(K-1)\sqrt{t}}\left(\frac{p}{(K-1)\sqrt{t}}\right)^{-1}\left\{1-\left(1-\frac{p}{(K-1)\sqrt{t}}\right)^{\rho_{t}}\right\}$
$=$ $1-\left(1-\frac{p}{(K-1)\sqrt{t}}\right)^{\rho_{t}}$
$\geq$ $\frac{p\rho_{t}}{(K-1)\sqrt{t}},$
where the last inequality uses $(1+x)^{n}\geq 1+nx$ for $x\geq-1$ and $n\in\mathbb{N}$. Then the expected sum of regret,
$\mathbb{E}\left[\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}=\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}\right]$ $\leq\sum_{t\in[T]\setminus\mathcal{E}_{T}}\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)$
$\leq\frac{2p\sqrt{T}}{K-1}\rho_{T}$
$=\frac{2p\sqrt{T}}{K-1}\frac{\log\frac{(T+1)^{2}}{\delta}}{\log(1/p)}.$
Thus,
$\mathbb{E}[\mathrm{Reg}(T)]\leq 6\delta\log T+\frac{32K^{2}}{(1-p)^{2}}\log\frac{2dT^{2}}{\delta}+\frac{2p\sqrt{T}}{K-1}\frac{\log\frac{(T+1)^{2}}{\delta}}{\log(1/p)}+8\sqrt{KT}\left(\frac{\sigma}{p}\sqrt{\log\frac{T+1}{\delta}}+1\right).\hfill\qed$
Appendix D Algorithm-agnostic Lower Bound of Regret Ignoring Unobserved Features
@@
Appendix E Technical Lemmas
@@
$\mathbb{P}\left(|{X_{T}-X_{0}}|\geq a\right)\leq 2\exp\left(-\frac{a^{2}}{2\sum_{t=1}^{T}\sigma_{t}^{2}}\right)$
@@
E.1 A Hoeffding bound for Matrices
@@
$\mathbb{P}\left(\left\|\sum_{\tau=1}^{t}\mathbf{M}_{\tau}-\mathbb{E}[{\mathbf{M}_{\tau}}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{\tau-1}}]\right\|_{2}\geq x\right)\leq 2d\exp\left(-\frac{x^{2}}{2tb^{2}}\right)$
@@
$\mathbb{P}\left(\left\|\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right\|_{2}\geq x\right)\leq\mathbb{P}\left(\lambda_{\max}\left(\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\geq x\right)+\mathbb{P}\left(\lambda_{\max}\left(-\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\geq x\right)$
@@
By Lieb’s theorem [Tropp, 2015] the mapping $\mathbf{D}\mapsto\exp(\mathbf{H}+\log\mathbf{D})$ is concave on positive symmetric matrices for any symmetric positive definite $H$. By Jensen’s inequality,
$\text{Tr}\mathbb{E}\left[\exp\left(v\sum_{\tau=1}^{t-1}\mathbf{D}_{\tau}+\log\exp(v\mathbf{D}_{t})\right)\right]\leq\text{Tr}\mathbb{E}\left[\exp\left(v\sum_{\tau=1}^{t-1}\mathbf{D}_{\tau}+\log\mathbb{E}\left[\left.\exp(v\mathbf{D}_{t})\right|\mathcal{F}_{t-1}\right]\right)\right]$
@@
E.2 A Bound for the Gram Matrix
@@
$\left\|\mathbf{I}_{K}-\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}\right\|_{2}\leq\epsilon,$
@@
$\leq\frac{2K}{1-p}\cdot\frac{1}{t},$
where the last inequality holds by Sherman-Morrison formula. By Hoeffding bound for matrix (Lemma 2), for $x>0$
@@
$\left\|\mathbf{I}_{K}-\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}\right\|_{2}\leq\epsilon,$
E.3 An error bound for the Lasso estimator
@@
The proof is similar to that of Lemma B.4 in [Kim et al., 2024], but we provide a new proof for the (unrestricted) minimum eigenvalue condition. Let $\mathbf{X}_{t}^{\top}:=(\mathbf{x}_{1},\ldots,\mathbf{x}_{t})\in[-1,1]^{d\times t}$ and $\mathbf{e}_{t}^{\top}:=(e_{1},\ldots,e_{t})\in\mathbb{R}^{t}$. We write $\mathbf{X}_{t}(j)$ and $\widehat{\mathbf{w}}_{t}(j)$ as the $j$-th column of $\mathbf{X}_{t}$ and $j$-th entry of $\widehat{\mathbf{w}}_{t}$, respectively. By definition of $\widehat{\mathbf{w}}_{t}$,
@@
$=$ $\|\bar{\mathbf{w}}\|_{1}-\sum_{i\in\bar{\mathcal{S}}}|{\widehat{\mathbf{w}}_{t}(i)-\bar{\mathbf{w}}(i)}|+\sum_{i\in[d]\setminus\bar{\mathcal{S}}}|{\widehat{\mathbf{w}}_{t}(i)}|$
@@
E.4 Eigenvalue bounds for the Gram matrix.

### v2

@@ E.5 Proof of Theorem 5
with probability at least $1-3\delta/t$, by Theorem 4. Summing over $t$ gives, (with probability at least $1-3\delta$),
@@
By Lemma 5 and the Hoeffding bound, with probability at least $1-\delta$
$\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}\neq\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}$
$\leq 2\left[\sum_{t\in[T]}\mathbb{I}\left(a_{t}\neq\widehat{a}_{t}\right)-\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)+\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)\right]$
$\leq 4\sqrt{2T\log\frac{2}{\delta}}+\frac{4\sqrt{T}}{K-1}+4\delta.$
Then the cumulative regret,
$\mathrm{Reg}(T)\leq\frac{32K^{2}}{(1-p)^{2}}\log\frac{2dT^{2}}{\delta}+4\sqrt{2T\log\frac{2}{\delta}}+\frac{4\sqrt{T}}{K-1}+4\delta+8\sqrt{KT}\left(\frac{\sigma}{p}\sqrt{\log\frac{t+1}{\delta}}+1\right),$
which completes the proof. ∎
Appendix F Algorithm-Agnostic Lower Bound of Regret Ignoring Unobserved Features
@@
Appendix G Technical Lemmas
@@
$\mathbb{P}\left(|{X_{T}-X_{0}}|\geq a\right)\leq 2\exp\left(-\frac{a^{2}}{2\sum_{t=1}^{T}\sigma_{t}^{2}}\right).$
@@
G.1 A Hoeffding bound for Matrices
@@
$\mathbb{P}\left(\left\|\sum_{\tau=1}^{t}\mathbf{M}_{\tau}-\mathbb{E}[{\mathbf{M}_{\tau}}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{\tau-1}}]\right\|_{2}\geq x\right)\leq 2d\exp\left(-\frac{x^{2}}{2tb^{2}}\right).$
@@
$\mathbb{P}\left(\left\|\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right\|_{2}\geq x\right)\leq\mathbb{P}\left(\lambda_{\max}\left(\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\geq x\right)+\mathbb{P}\left(\lambda_{\max}\left(-\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\geq x\right).$
@@
By Lieb’s theorem (Tropp, 2015) the mapping $\mathbf{D}\mapsto\exp(\mathbf{H}+\log\mathbf{D})$ is concave on positive symmetric matrices for any symmetric positive definite $H$. By Jensen’s inequality,
$\text{Tr}\mathbb{E}\left[\exp\left(v\sum_{\tau=1}^{t-1}\mathbf{D}_{\tau}+\log\exp(v\mathbf{D}_{t})\right)\right]\leq\text{Tr}\mathbb{E}\left[\exp\left(v\sum_{\tau=1}^{t-1}\mathbf{D}_{\tau}+\log\mathbb{E}\left[\left.\exp(v\mathbf{D}_{t})\right|\mathcal{F}_{t-1}\right]\right)\right].$ (24)
@@
G.2 A Bound for the Gram Matrix
@@
$\left\|\mathbf{I}_{K}-\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}\right\|_{2}\leq\epsilon.$
@@
$\leq\frac{2K}{1-p}\cdot\frac{1}{t}.$ (25)
Note that the last inequality (Eq. 25) holds by Sherman-Morrison formula. By Hoeffding bound for matrix (Lemma 2), for $x>0$, it holds that:
@@
$\left\|\mathbf{I}_{K}-\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}\right\|_{2}\leq\epsilon.$
G.3 An error bound for the Lasso estimator
@@
The proof is similar to that of Lemma B.4 in (Kim et al., 2024), but we provide a new proof for the (unrestricted) minimum eigenvalue condition. Let $\mathbf{X}_{t}^{\top}:=(\mathbf{x}_{1},\ldots,\mathbf{x}_{t})\in[-1,1]^{d\times t}$ and $\mathbf{e}_{t}^{\top}:=(e_{1},\ldots,e_{t})\in\mathbb{R}^{t}$. We write $\mathbf{X}_{t}(j)$ and $\widehat{\mathbf{w}}_{t}(j)$ as the $j$-th column of $\mathbf{X}_{t}$ and $j$-th entry of $\widehat{\mathbf{w}}_{t}$, respectively. By definition of $\widehat{\mathbf{w}}_{t}$,
@@
$=$ $\|\bar{\mathbf{w}}\|_{1}-\sum_{i\in\bar{\mathcal{S}}}|{\widehat{\mathbf{w}}_{t}(i)-\bar{\mathbf{w}}(i)}|+\sum_{i\in[d]\setminus\bar{\mathcal{S}}}|{\widehat{\mathbf{w}}_{t}(i)}|,$
@@
G.4 Eigenvalue bounds for the Gram matrix.
@@
G.5 A Bound for the Probability of Exploration
Lemma 5.
For each $t$, let $\widehat{a}_{t}=\mathop{\mathrm{argmax}}_{a\in[K]}\widetilde{\mathbf{x}}_{a}^{\top}\widehat{\boldsymbol{\mu}}_{t}$ denote the maximizing action based on the estimator $\widehat{\boldsymbol{\mu}}_{t}$. Then the action $a_{t}$ chosen by the resampling scheme in Algorithm 1 and Algorithm 2 satisfies,
$\sum_{t=1}^{T}\mathbb{P}(a_{t}\neq\widehat{a}_{t})\leq\frac{2\sqrt{T}}{(K-1)}+2\delta.$
Proof.
Since the algorithm resampling at most $\rho_{t}$ times, for a fixed $t\in[T]$,
$\mathbb{P}(a_{t}\neq\hat{a}_{t})$ =($=\mathbb{P}\left(\{\widetilde{a}_{t}=a_{t}\}\cap\{a_{t}\neq\hat{a}_{t}\}\right)+\mathbb{P}\left(\{\widetilde{a}_{t}\neq a_{t}\}\cap\{a_{t}\neq\hat{a}_{t}\}\right)$
($\leq\mathbb{P}\left(\{\widetilde{a}_{t}=a_{t}\}\cap\{a_{t}\neq\hat{a}_{t}\}\right)+\mathbb{P}(\widetilde{a}_{t}\neq a_{t})$
=($=\mathbb{P}\left(\{\widetilde{a}_{t}=a_{t}\}\cap\{a_{t}=k\}\right)+\underbrace{\mathbb{P}(\widetilde{a}_{t}\neq a_{t})}_{\text{Failure of resampling}}\quad\text{for }k\neq\hat{a}_{t},$ (27)
where $\mathbb{P}(a_{t}=k)=t^{-1/2}/(K-1)$ and $\mathbb{P}(\widetilde{a}_{t}=a_{t})=p$, defined by Algorithm 1 and Eq. 9, respectively. For the first term in Eq. 27, by union bound it holds that:
($\mathbb{P}\left(\{\widetilde{a}_{t}=a_{t}\}\cap\{a_{t}=k\}\right)$ $=\mathbb{P}\left(\bigcup_{m=1}^{\rho_{t}}\{\text{Resampling success at trial }m\}\cap\{a_{t}=k\}\right)$
$\leq\sum_{m=1}^{\rho_{t}}\mathbb{P}\left(\{\text{Resampling success at trial }m\}\cap\{a_{t}=k\}\right),$
$\leq\mathbb{P}(a_{t}=k)$
$=\frac{1}{\sqrt{t}(K-1)}.$
which, combined with Eq. 27, results in:
$\mathbb{P}(a_{t}\neq\hat{a}_{t})\leq\frac{1}{\sqrt{t}(K-1)}+\mathbb{P}(\text{Resampling failure})$
By the definition of $\rho_{t}$, the probability that the resampling fails is bounded by $\delta/(t+1)^{2}$. Thus, the probability of the event $\{a_{t}\neq\widehat{a}_{t}\}$ is
$\mathbb{P}(a_{t}\neq\hat{a}_{t})\leq\frac{1}{\sqrt{t}(K-1)}+\frac{\delta}{(t+1)^{2}}$
Summing up over $t\in[T]$ completes the proof. ∎